OpenStudy (anonymous):

Keisha's parents want to save twenty thousand dollars in her college savings account over the next fifteen years. They have eight thousand dollars to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth. A)two point five percent B)two point seven percent C)ten percent D)twelve percent

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Principal P = 8000 Number of years n = 15 Let interest rate be r percent per year Final Amount A = 20000 The formula to use is: A = P( 1 + r * t ) Plug in the values: 20000 = 8000(1 + 15r) Solve for r.

OpenStudy (ranga):

what did you get for r? If you need further help in solving for r just let me know.